Q:- The point on the circle 16(x2+y2)+48x-8y-43=0 nearest to the line 8x-4y+73=0 is _______________

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear tapasranjan

16(x2+y2)+48x-8y-43=0

 x2+y2+3x-1/2 y- 43/16=0

 (x+3/2)2 + (y-1/4)2 =5

now find the equation of line perpendicular to the given line and passes through center of the circle.

 it will come  x+2y+1 =0

now find the point of intersection of this line to the given circle.

points will come (-7/2 ,5/4)  and (1/2,-3/4)

now find the length of perpendicual from these point to the given line . point which correspond to small length of pendicular is the desired point

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin